OpenStudy (anonymous):

the temperature, T, in degrees celsius, of a yam put into a 200 degrees C oven is given as a function of time, t, in minutes, by: T = a(1-e^-kt)+ b. a)if the yam starts at 20 degrees C, find a and b. b)if the temperature of the yam is initially increasing at 2 degrees C per minute, find k.

OpenStudy (solomonzelman):

Physics is a correct section for this question, and there is a link for you. http://openstudy.com/study#/groups/Physics

OpenStudy (anonymous):

thanks...

zepdrix (zepdrix):

\[\Large\bf T(t)\quad=\quad a(1-e^{kt})+b\] `a) If the yam starts at 20 degrees C, find a and b.` So the way to interpret this is, at time t=0 the temperature T of the yam is 20. \(\Large\bf T(0)=20\)

zepdrix (zepdrix):

\[\Large\bf 20\quad=\quad a(1-e^{0k})+b\]Understand how I set that up? I set it equal to a `temperature` of 20, and then plugged in 0 for the `time`.

OpenStudy (anonymous):

okay i understand that so far. so now we solve for a? would it be a = (20-b) / (1-e^0k) ?

zepdrix (zepdrix):

No it looks like we'll start by solving for b. the exponential simplifies down a bit, \(\Large \bf e^{0k}=1\)

zepdrix (zepdrix):

So our expression is:\[\Large\bf 20\quad=\quad a(1-1)+b\]

OpenStudy (anonymous):

so b = 20. then plug it back into the equation?

zepdrix (zepdrix):

yah let's do that and see if we can figure out a.. hmm

OpenStudy (anonymous):

I'm confused! ): how do you find a if a(0) = 0?

zepdrix (zepdrix):

We'll have to find a `without` plugging in the T(20)=0. We can safely plug in b=20 though.

OpenStudy (anonymous):

so you mean 20 = a(1-e^kt) + (20) ?

zepdrix (zepdrix):

\[\Large\bf T(t)\quad=\quad a(1-e^{kt})+b\] We used T(20)=0 to find b. b=20. So now we plug that into our function.\[\Large\bf T(t)\quad=\quad a(1-e^{kt})+20\]

OpenStudy (anonymous):

but now how do we solve for a if we are missing t?

OpenStudy (anonymous):

wait. I'm sorry. i just realized that i typed the wrong equation. it should be T= a(1 - e^-kt) + b does that change things?

zepdrix (zepdrix):

\[\Large\bf T(t)\quad=\quad a(1-e^{-kt})+20\]No, luckily it doesn't change anything we've done so far.

zepdrix (zepdrix):

`The yam is put into a 200 degree C oven,` Hmm I'm trying to figure out how this plays into our problem :( It probably helps us figure out `a` somehow. hmmm

zepdrix (zepdrix):

I'm not sure why Solomon tried to kick you over to the Physics section :P lol This is as much a math problem as it is physics XD

OpenStudy (anonymous):

yea, thats what i thought lol.

zepdrix (zepdrix):

So the `initial` difference in temperature between the yam and the oven is 200 - 20 = 180 degrees C. I think we need to use this somehow.. hmmm thinkingggggggggg

zepdrix (zepdrix):

@hartnn @Zarkon @Loser66

OpenStudy (anonymous):

D: I'm sorry i have no idea what to do next

zepdrix (zepdrix):

ya I'm a little stumped here :d grr

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

@satellite73

zepdrix (zepdrix):

the @ paging thing isn't working right now :( grr

hartnn (hartnn):

`a yam put into a 200 degrees C oven` means the max value the yam can achieve in the oven is = 200 degreesC T = a(1-e^-kt)+ b. will be maximum when you do not subtract anything from 1 that is, when e^-kt = 0 so, Tmax = 200 = a(1-0)+b from this you can find 'a' :) see if it makes sense.

hartnn (hartnn):

hint for part b) ` increasing at 2 degrees C per minute` ----> dT/dt = +2

OpenStudy (anonymous):

okay so i got 1 = 180.

OpenStudy (anonymous):

now for b. i know you plug in 2 for dT/dt and find the derivative but I'm kind of confused about it :\

hartnn (hartnn):

so you know you have to find the derivative, did you try finding it ? what did u get ?